OpenStudy (anonymous):

how to simplify a radical expression. ex: Simplify the prob 3Sqrrt(-243x^3y^10) i already know nSqrrt(a) * nSqrrt(B) = nSqrrt(AB)

OpenStudy (anonymous):

ok so \(\sqrt[3]{x^3}=x\) for a start

OpenStudy (anonymous):

and since \(243=3^5\) we have \(\sqrt[3]{3^5}=3\sqrt[3]{3^2}\)

OpenStudy (anonymous):

finally \(\sqrt[3]{y^{10}}=3\sqrt[3]{y}\)